# Upper bounds on the highest phonon frequency and superconducting temperature from fundamental physical constants

### v1

@@ 3.2 Other coupling mechanisms
The above discussion also applies to systems where superconductivity is related to mechanisms other than the electron-phonon coupling. For electron-electron pairing, $T_{c}$ depends on the Fermi energy $E_{\rm F}$ and is on the order of about $\frac{E_{\rm F}}{10^{2}}$ Varma (2012); Uemura et al. (1991). $E_{\rm F}\propto\frac{\hbar^{2}}{m_{e}a_{\rm B}^{2}}\propto\frac{m_{e}e^{4}}{\hbar^{2}}$ and depends on fundamental constants as $E_{\rm R}$ in Eq. (2) does. Hence, $T_{c}$ in the electron-electron systems is also governed by fundamental constants as it is in the electron-phonon superconductors.
Assuming that the upper limit to $E_{\rm F}$ is set by $E_{\rm R}$ in Eq. (2) and recalling that $T_{c}$ in the electron-electron systems is on the order of $\frac{E_{\rm F}}{10^{2}}$ Varma (2012); Uemura et al. (1991), we note that the upper bound to $T_{c}$ in these systems is close to $T_{c}^{u}$ since $C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}$ in Eq. (8) with $C\approx 0.1$ is on the order of about $\frac{1}{10^{2}}$.
@@ 3.3 Fundamental constants and search for room-temperature superconductivity
Having seen that the overall scale of $T_{c}^{u}$ is set by fundamental physical constants, we now ask the important question of what would be the effects of fundamental constants having different values. This question is intensely researched in high-energy physics and astrophysics and is related to fine-tuning of the Universe and associated grand challenges in modern science Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003); Barrow and Webb (2006). How would superconductivity change were fundamental constants (e.g., $\hbar$, $m_{e}$, $e$ and $\frac{m_{e}}{m_{p}}$ in Eq. (8)) to take different values? This question is well-posed because it is possible to change $T_{c}^{u}$ in Eq. (8) by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta=\frac{m_{e}}{m_{p}}$ and $\alpha=\frac{e^{2}}{\hbar c}$ unchanged and hence maintain the rest of essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact Barrow (2003). Were different fundamental constants to give $T_{c}^{u}$ on the order of, for example, $10^{-6}$ K or lower, superconductivity would be unobserved. Were different fundamental constants to give $T_{c}^{u}$ on the order $10^{6}$ K, many superconductors would have $T_{c}$ in excess of 300 K. Consistent with Eqs. (8)-(9) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to currently observed fundamental constants, we observe superconductivity in the temperature range $T\lesssim$ 100-200 K. This understandably stimulates the current research into finding systems with $T_{c}$ of 300 K and above. Therefore, the very existence of the current line of enquiry to identify systems with $T_{c}$ above 300 K is itself due to the values of fundamental constants currently observed.
@@ 4 Fundamental constants and observability of phenomena in condensed matter physics
The discussion in the previous section implies that fundamental constants can affect observability of an entire condensed matter phenomenon such as superconductivity. This implication is wider than what was discussed in earlier work Trachenko (2023a); Trachenko and Brazhkin (2020); Trachenko (2023b); Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018); Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Trachenko et al. (2021); Brazhkin (2023); Trachenko et al. (2020); Trachenko (2024, 2023c) where fundamental constants were shown to set a bound on a property (e.g., viscosity, thermal conductivity, diffusivity, or speed of sound) which is already observed.
@@
The observability of superconductivity as a result of current values of fundamental constants can be extended more generally to the observability of other phenomena and phase transitions. Indeed, different fundamental constants substantially increasing $\omega_{\rm m}^{\rm u}$ in Eq. (4) would result in $T\ll\hbar\omega_{\rm m}$ and hence $T\ll\hbar\omega_{\rm D}$ ($\omega_{\rm D}$ is Debye frequency) at typical planetary temperatures. This would result in the solid properties being always quantum and the classical regime $T>\hbar\omega_{\rm D}$ unobservable. This would also suppress or arrest second-order structural phase transitions because of the smallness of the phonon entropy when $T\ll\hbar\omega_{\rm D}$ Landau and Lifshitz (1970).
@@
Another example is the first-order melting transition where the characteristic slopes of the pressure-temperature melting lines are set by fundamental constants Trachenko (2024). This implies that different values of these constants would result in melting either (a) always taking place at typical planetary conditions and no solids existing or (b) being unobservable (in this case, the ice melting temperature would be above typical planetary temperatures implying no water-based life).
@@
Linking the discussion the previous paragraph and in Section 3.3, we note that current fundamental constants enable diverse phenomena such as water-based life and superconductivity.
@@
This and similar future discussion deepen our understanding of how fundamental constants affect observability of entire new phenomena at condensed matter length and energy scales. This helps to fill the large gap Trachenko (2023b) between the scales involved in nuclear physics and astrophysics where fundamental constants were shown to give rise to different key effects such as the stability of nuclei, star formation, and heavy nuclei synthesis Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). Observability of each effect and phenomenon imposes constraints on fundamental constants Barrow (2003). Discussing different constraints along a continuous path of length and energy scales, including condensed matter scales, is likely to bring about new insights into understanding fundamental constants.
@@ Acknowledgements.
In our density functional theory calculations, we use the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional Perdew et al. (1996), an energy cutoff of $1200$ eV and a $\mathbf{k}$-point grid of spacing $2\pi\times 0.025$-1 to sample the electronic Brillouin zone. We relax cell parameters and internal coordinates to obtain a pressure to within $10^{-4}$ GPa of the target pressure and forces smaller than $10^{-5}$ eV/Å. We calculate the phonon dispersion using the finite displacement method Kunc and Martin (1978) in conjunction with nondiagonal supercells Lloyd-Williams and Monserrat (2015) with a $4\times 4\times 4$ coarse $\mathbf{q}$-point grid to sample the vibrational Brillouin zone, and we use Fourier interpolation to calculate the phonon frequencies along a high symmetry path in the Brillouin zone.
@@
The critical temperature is obtained by solution of the Eliashberg equations Marsiglio (2020), whose only inputs are the Eliashberg function and a value for the Coulomb pseudopotential $\mu^{*}$. We can therefore write $T_{c}=T_{c}[\alpha^{2}F](\mu^{*})$, a functional of $\alpha^{2}F$ and a function of $\mu^{*}$.
@@
In particular, noting that Eq. 11 is linear in $\alpha^{2}F$, we can satisfy Eq. 12 by taking
@@
where mapping $f\rightarrow f^{2}$ ensures positivity (Eq. 13) and $G$ is an envelope function, which we take as
@@
so that the limits in Eqs. 14 and 15 hold, and are approached smoothly. We take the width of the envelope to be $\sigma=300$ K, but this value makes little difference to the resulting $T_{c}$.
@@
by representing $f$ on a uniform grid of frequency points and employing the BFGS algorithm Fletcher (1987). We begin the optimization with an initial guess of $f=1$.
merlin.mbs apsrev4-1.bst 2010-07-25 4.21a (PWD, AO, DPC) hacked

### v2

@@ 3.2 Other coupling mechanisms
The above discussion also applies to systems where superconductivity is related to mechanisms other than the electron-phonon coupling. For electron-electron pairing, $T_{c}$ depends on the Fermi energy $E_{\rm F}$ and is on the order of about $\frac{E_{\rm F}}{10^{2}}$ [60, 75]. $E_{\rm F}\propto\frac{\hbar^{2}}{m_{e}a_{\rm B}^{2}}\propto\frac{m_{e}e^{4}}{\hbar^{2}}$ and depends on fundamental constants as $E_{\rm R}$ in Eq. (2) does. Hence, $T_{c}$ in the electron-electron systems such as cuprates also depends on on fundamental constants as it is in the electron-phonon superconductors.
Noting that the upper limit to $E_{\rm F}$ is set by $E_{\rm R}$ in Eq. (2) and recalling that $T_{c}$ in the electron-electron systems is on the order of $\frac{E_{\rm F}}{10^{2}}$ [60, 75], we observe that the upper bound to $T_{c}$ in these systems is close to $T_{c}^{u}$ because $C\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}$ in Eq. (8) with $C\approx 0.1$ is on the order of about $\frac{1}{10^{2}}$.
@@ 3.3 Fundamental constants and search for room-temperature superconductivity
Having seen that the overall scale of $T_{c}^{u}$ is set by fundamental physical constants, we now ask the important question of what would be the effects of fundamental constants having different values. This question is intensely researched in high-energy physics and astrophysics and is related to fine-tuning of the Universe and associated grand challenges in modern science [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. How would superconductivity change were fundamental constants (e.g., $\hbar$, $m_{e}$, $e$ and $\frac{m_{e}}{m_{p}}$ in Eq. (8)) to take different values? This question is well-posed because it is possible to change $T_{c}^{u}$ in Eq. (8) by varying $\hbar$, $m_{e}$, $e$ and $m_{p}$ while keeping $\beta=\frac{m_{e}}{m_{p}}$ and $\alpha=\frac{e^{2}}{\hbar c}$ unchanged and hence maintain the rest of essential processes (e.g. those involved in stellar formation and evolution, synthesis of heavy elements, stability of ordered structures and so on) intact [1].
We observe that were different fundamental constants to give $T_{c}^{u}$ on the order of, for example, $10^{-6}$ K or lower, superconductivity would be unobserved. Were different fundamental constants to give $T_{c}^{u}$ on the order $10^{6}$ K, many superconductors would have $T_{c}$ in excess of 300 K. Consistent with Eqs. (8)-(9) giving $T_{c}^{u}$ on the order of $10^{2}-10^{3}$ K due to currently observed fundamental constants, we observe superconductivity in the temperature range $T\lesssim$ 100-200 K. This understandably stimulates the current research into finding systems with $T_{c}$ of 300 K and above.
Therefore, the very existence of the current line of enquiry to identify systems with $T_{c}$ above 300 K is itself due to the values of fundamental constants currently observed.
@@ 4 Fundamental constants and observability of phenomena in condensed matter physics
The discussion in the previous section implies that fundamental constants can affect observability of an entire condensed matter phenomenon such as superconductivity. This implication is wider than what was discussed in earlier work [15, 16, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13, 31, 12, 14] where fundamental constants were shown to set a bound on a property (e.g., viscosity, thermal conductivity, diffusivity, or speed of sound) which is already observed.
@@
The observability of superconductivity as a result of current values of fundamental constants can be extended more generally to the observability of other phenomena and phase transitions. Indeed, different fundamental constants substantially increasing $\omega_{\rm m}^{\rm u}$ in Eq. (4) would result in $T\ll\hbar\omega_{\rm m}$ and hence $T\ll\hbar\omega_{\rm D}$ ($\omega_{\rm D}$ is Debye frequency) at typical planetary temperatures. This would result in the solid properties being always quantum and the classical regime $T>\hbar\omega_{\rm D}$ unobservable. This would also suppress or arrest second-order structural phase transitions because of the smallness of the phonon entropy $S\propto\left(\frac{T}{\hbar\omega_{\rm D}}\right)^{3}$ when $T\ll\hbar\omega_{\rm D}$ [76].
@@
Another example is the first-order melting transition where melting temperature and characteristic slopes of the pressure-temperature melting lines are set by fundamental constants [30]. This implies that different values of these constants would result in melting either (a) always taking place at typical planetary conditions and no solids existing or (b) being unobservable (in this case, the ice melting temperature would be above typical planetary temperatures implying no water-based life).
@@
Linking the discussion in the previous paragraph and in Section 3.3, we note that current fundamental constants enable diverse phenomena such as water-based life and superconductivity.
@@
This and similar future discussion deepen our understanding of how fundamental constants affect observability of entire new phenomena at condensed matter length and energy scales. This helps to fill the large gap [14] between the scales involved in nuclear physics and astrophysics where fundamental constants were shown to give rise to different key effects such as the stability of nuclei, star formation, and heavy nuclei synthesis [1, 2, 3, 4, 5, 6, 7, 8, 9]. Observability of each effect and phenomenon imposes constraints on fundamental constants [1]. Discussing different constraints along a continuous path of length and energy scales, including condensed matter scales, is likely to bring about new insights into understanding fundamental constants.
There may be a sense in which bounds to different properties and observability of phenomena can be expected to be related to fundamental constants on general grounds. We add two remarks in this regard. First, it is important to show how and why this happens. Indeed, one view holds that the essence of a physical theory is to provide a relationship between different physical properties and experimental outcomes [77]. Our current discussion provides such relationships, including the relationship between bounds on vibration frequency, $T_{c}$, observability of phase transitions on one hand and fundamental constants on the other hand. Second, it is not clear apriori whether bounds can be expressed solely in terms of fundamental constants: recently discussed bounds depend on the combination of fundamental constants and external parameters such as temperature [18, 21, 22, 23, 24, 25, 26]. On the other hand, our bound on oscillation frequency (4) solely depends on fundamental constants. This is a more general and interesting result. This is particularly so in view that phenomena such as superconductivity, phase transitions and so on involve many-body collective effects which are considered as emergent and not reducible to fundamental constants (see Introduction).
@@ Acknowledgements.
In our density functional theory calculations, we use the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [78], an energy cutoff of $1200$ eV and a $\mathbf{k}$-point grid of spacing $2\pi\times 0.025$-1 to sample the electronic Brillouin zone. We relax cell parameters and internal coordinates to obtain a pressure to within $10^{-4}$ GPa of the target pressure and forces smaller than $10^{-5}$ eV/Å. We calculate the phonon dispersion using the finite displacement method [79] in conjunction with nondiagonal supercells [80] with a $4\times 4\times 4$ coarse $\mathbf{q}$-point grid to sample the vibrational Brillouin zone, and we use Fourier interpolation to calculate the phonon frequencies along a high symmetry path in the Brillouin zone.
@@
The critical temperature is obtained by solution of the Eliashberg equations [72], whose only inputs are the Eliashberg function and a value for the Coulomb pseudopotential $\mu^{*}$. We can therefore write $T_{c}=T_{c}[\alpha^{2}F](\mu^{*})$, a functional of $\alpha^{2}F$ and a function of $\mu^{*}$.
@@
In particular, noting that Eq. (11) is linear in $\alpha^{2}F$, we can satisfy Eq. (12) by taking
@@
where mapping $f\rightarrow f^{2}$ ensures positivity (Eq. (13)) and $G$ is an envelope function, which we take as
@@
so that the limits in Eqs. (14) and (15) hold, and are approached smoothly. We take the width of the envelope to be $\sigma=300$ K, but this value makes little difference to the resulting $T_{c}$.
@@
by representing $f$ on a uniform grid of frequency points and employing the BFGS algorithm [81]. We begin the optimization with an initial guess of $f=1$.
A code to perform this optimization is available [82].